OpenStudy (therealmeeeee):

Simplify: 4x5 + 2x5 6 6x5 6x 6x10

OpenStudy (arthur_ser):

4x5 + 2x5=5x(2+4)

OpenStudy (the_fizicx99):

B 4x^5 + 2x^5 = 6x^5

OpenStudy (therealmeeeee):

Simplify: y7 • y3 y9 y4 y10 y21

OpenStudy (the_fizicx99):

y^10

OpenStudy (therealmeeeee):

Simplify: 34 9 81 16 8

OpenStudy (the_fizicx99):

34?

OpenStudy (the_fizicx99):

Simplify what..

OpenStudy (therealmeeeee):

Simplify: 34 9 81 16 8

OpenStudy (therealmeeeee):

yeah 34

OpenStudy (therealmeeeee):

no not 34

OpenStudy (the_fizicx99):

Idk what to simplify with.

OpenStudy (therealmeeeee):

3 than a lil 4

OpenStudy (therealmeeeee):



OpenStudy (the_fizicx99):

3^4 is the same as 3 * 3 * 3 * 3 = 81 3*3 = 9 *3 = 27 * 3 = 81

OpenStudy (therealmeeeee):

Simplify: x7 x3 x2 x4 x7 - x3 x10

OpenStudy (the_fizicx99):

x7/x3 ? Divide?

OpenStudy (therealmeeeee):

no *

OpenStudy (the_fizicx99):

Are you adding, subtracting, multiplying, dividing?

OpenStudy (therealmeeeee):

multipling

OpenStudy (the_fizicx99):

Then add them 7 + 3 = x^?

OpenStudy (the_fizicx99):

Add the exponents.. 7+3 = ??

OpenStudy (the_fizicx99):

@TheRealMeeeee ?